OpenStudy (anonymous):

how to convert -16^1/2 to radical expression

Parth (parthkohli):

\[a^{1/2} = \sqrt{a}\]

OpenStudy (anonymous):

what are the actual steps to convert -16^1/2 to a radical expression

Parth (parthkohli):

Just use the property that I mentioned.

OpenStudy (anonymous):

\[\sqrt[2]{16}\] is this how I would write it

OpenStudy (amistre64):

yes